OpenStudy (seratul):

Use the diagonals to tell whether a parallelogram with the given vertices is a rectangle, rhombus, or square. Give all names that apply. Q(-8,-2) T(-6,8) W(4,6) Z(2,-4)

OpenStudy (vespersaviour):

Search up images of graph paper then try graphing these @Seratul

OpenStudy (seratul):

I am not allowed to graph them. It would be too easy that way o.O

OpenStudy (vespersaviour):

Can you guess from this graph?

OpenStudy (seratul):

Square and rectangle?

OpenStudy (vespersaviour):

I think you have square but look at this one and see. Im not to sure.. Im only good at graphing points XD

OpenStudy (seratul):

Then I don't think you can help me xD. I need to make sure it is a certain shape with diagonals.

OpenStudy (vespersaviour):

Eh well i tried XD Sowwie

OpenStudy (vishweshshrimali5):

Here are some points regarding diagonals of each of the mentioned figures: 1. Rectangle -> Diagonals are equal and bisect each other but not perpendicularly 2. Rhombus -> Diagonals are perpendicular bisectors of each other but not equal 3. Square -> Diagonals equal and perpendicular bisectors of each other 4. Parallelogram -> Diagonals bisect each other.. no need to be equal

OpenStudy (vishweshshrimali5):

Any doubt with this?

OpenStudy (seratul):

I'm pretty sure the diagonals of rectangles bisect each other perpendicularly.

OpenStudy (seratul):

Anyways, I still think it would be a square and rectangle.

OpenStudy (vishweshshrimali5):

A rhombus is a rectangle with all sides equal or diagonals perpendicular bisecting each other

OpenStudy (seratul):

So all three?

OpenStudy (vishweshshrimali5):

Q(-8,-2) T(-6,8) W(4,6) Z(2,-4) What is the order of the vertices ?

OpenStudy (vishweshshrimali5):

I mean the figures should be QTWZ or QWTZ or anything else?

OpenStudy (seratul):

I think the basic shape is what the previous person drew above.

OpenStudy (vishweshshrimali5):

Lite.. it is QTWZ...

OpenStudy (vishweshshrimali5):

So anyways... start by calculating the slope of QW and TZ

OpenStudy (vishweshshrimali5):

Can you do this?

OpenStudy (seratul):

Yea. QW: 6--2/4--8 = 8/12 = 2/3 TZ: -4-8/2--6 = -12/8= -3/2 They are negative reciprocals so they are parallel.

OpenStudy (seratul):

But we know that since they are parallelograms which is given.

OpenStudy (vishweshshrimali5):

Naah naah!! Because the slopes are negative reciprocals they are perpendicular... for parallel the slopes must be the same

OpenStudy (seratul):

Whoops, my bad xD.

OpenStudy (vishweshshrimali5):

Thus, diagonals QW and TZ are perpendicular.. thus, it can either be a square of a rhombus.. now you just need to see whether diagonals are equal or not

OpenStudy (vishweshshrimali5):

Use the distance formula for finding the length of QW and TZ

OpenStudy (seratul):

Why can't it be a rectangle?

OpenStudy (seratul):

Oh wait, the diagonals aren't perpendicular.

OpenStudy (vishweshshrimali5):

Because for a rectangle the diagonals are not perpendicular to each other...

OpenStudy (seratul):

Okay, so know we use midpoint formula?

OpenStudy (vishweshshrimali5):

No.. we use distance formula

OpenStudy (vishweshshrimali5):

We want to find out the length of QW and TZ

OpenStudy (seratul):

QW: (4--8)^2 + (6--2)^2 = 12^2 + 8^2 = 144 +64 = sqrt 208 = 4 radical 13

OpenStudy (seratul):

TZ: (2--6)^2 + (-4-8)^2 = 8^2 + -12^2 = 64 + 144 + sqrt 208 = 4 radical 13

OpenStudy (seratul):

Okay, they are the same length.

OpenStudy (vishweshshrimali5):

Great!

OpenStudy (vishweshshrimali5):

So the figure is a square

OpenStudy (seratul):

Okay, i'm going to have to memorize the properties of the parallelograms. Thank you very much!

OpenStudy (vishweshshrimali5):

:) Your welcome